how many ways can 3 kids be selected from 12?
Answer by jim_thompson5910(35256)

There are 12 ways to choose the first kid

There are 12-1 = 11 ways to choose the next kid (assuming the first kid can't be placed back in the mix and potentially chosen again)

There are 12-2 = 10 ways to choose the third kid.

So there are 12*11*10 = 1320 different ways to choose 3 kids where order matters.

If order does not matter, then you'll have to divide by 6 because there are 3! = 6 ways to arrange any particular group. So if order does not matter, then there are 1320/6 = 220 different ways.

Summary:

If order does matter, then there are 1320 different ways

If order does NOT matter, then there are 220 different ways



--------------------------------------------------------------------------------------------------------------

If you need more help, you can email me at jim_thompson5910@hotmail.com

To find other ways to contact me, you can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------